# Retrieval-Augmented Generation for Reliable Interpretation of Radio Regulations

### v1

@@ Abstract
We study question answering in the domain of radio regulations, a legally sensitive and high-stakes area. We propose a telecom-specific Retrieval-Augmented Generation (RAG) pipeline and introduce, to our knowledge, the first multiple-choice evaluation set for this domain, constructed from authoritative sources using automated filtering and human validation. To assess retrieval quality, we define a domain-specific retrieval metric, under which our retriever achieves approximately 97% accuracy. Beyond retrieval, our approach consistently improves generation accuracy across all tested models. In particular, while naïvely inserting documents without structured retrieval yields only marginal gains for GPT-4o (less than 1%), applying our pipeline results in nearly a 12% relative improvement. These findings demonstrate that carefully targeted grounding provides a simple yet strong baseline and an effective domain-specific solution for regulatory question answering. All code and evaluation scripts, along with our derived question–answer dataset, are available at https://github.com/Zakaria010/Radio-RAG.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language processing, achieving state-of-the-art performance in summarization, translation, and question answering. However, despite their versatility, LLMs are prone to generating false or misleading content, a phenomenon commonly referred to as hallucination Huang_2025 ; sahoo2024comprehensivesurveyhallucinationlarge . While sometimes harmless in casual applications, such inaccuracies pose significant risks in domains that demand strict factual correctness, including medicine, law, and telecommunications. In these settings, misinformation can have severe consequences, ranging from financial losses to safety hazards and legal disputes.
@@
To address these challenges, we introduce a domain-specialized Retrieval-Augmented Generation (RAG) pipeline for regulatory question answering. Our system leverages authoritative external resources to ground LLM outputs, thereby reducing hallucinations and improving reliability lewis2021retrieval ; gupta2024comprehensive . To evaluate this approach, we construct a dedicated dataset of regulatory questions and answers derived from the ITU Radio Regulations ITURadioRegulations2024online , validated through both automated checks and human review. Our contributions are threefold:
@@
    We present an extensive empirical evaluation across multiple LLMs, showing that our pipeline consistently improves accuracy. For instance, on GPT-4o we achieve a +11.9% absolute accuracy improvement over vanilla prompting, while retrieval accuracy alone reaches up to 97%.
@@ 2.1 LLMs
LLMs are transformer-based neural networks vaswani2023attentionneed ; xiao2025foundationslargelanguagemodels trained on massive text corpora to acquire broad language understanding and reasoning capabilities. Their performance can be further enhanced through several complementary strategies. Supervised fine-tuning (SFT) adapts pretrained models to specific domains or tasks by updating their parameters on curated labeled data raffel2023exploringlimitstransferlearning ; hu2021loralowrankadaptationlarge ; xu2023parameterefficientfinetuningmethodspretrained , while reinforcement learning from human feedback (RLHF) aligns their behavior with human preferences through reward-based optimization ouyang2022traininglanguagemodelsfollow . Both methods directly modify the model’s parametric memory, embedding new knowledge into its weights, yet this knowledge remains static and can quickly become obsolete, especially in rapidly evolving or highly regulated domains ovadia2024finetuningretrievalcomparingknowledge .
In contrast, prompt learning zhou2022large ; kharrat2024acing and prompt engineering sahoo2024systematic guide model behavior without altering its parameters, using carefully crafted instructions or few-shot exemplars to induce desired task behavior at inference time. Relatedly, chain-of-thought (CoT) wei2022chain prompting encourages models to generate intermediate reasoning steps before producing an answer, which improves multi-step reasoning but does not expand the model’s knowledge base. While these approaches enhance task adaptability and reasoning ability, they cannot update or extend the underlying knowledge stored in the model’s parameters, leaving LLMs vulnerable to hallucinations, especially in specialized or fast-changing domains.
@@ 3 Related Works
The telecommunications domain poses distinctive challenges for language models due to its dense technical standards, highly structured documents, and precise terminology requirements. Recent efforts have begun adapting Retrieval-Augmented Generation (RAG) frameworks to this space. Bornea et al. bornea2024telcorag introduce Telco-RAG, a system designed to process 3GPP specifications, while Saraiva et al. saraiva2024telcodpr propose Telco-DPR, which evaluates retrieval models on a hybrid dataset combining textual and tabular inputs from telecom standards. Maatouk et al. maatouk2023teleqna contribute TeleQnA, a benchmark for assessing LLM knowledge of telecommunications. Subsequently Tele-LLMs maatouk2025telellmsseriesspecializedlarge were presented as a family of domain-specialized models trained on curated datasets (Tele-Data) and evaluated on a large-scale benchmark (Tele-Eval), demonstrating that targeted pretraining and parameter-efficient adaptation yield substantial improvements over general-purpose LLMs. More recently, Zou et al. zou2024telecomgpt developed TelecomGPT, a telecom-specific LLM built via continual pretraining, instruction tuning, and alignment tuning, and evaluated on new benchmarks such as Telecom Math Modeling, Telecom Open QnA, and Telecom Code Tasks; TelecomGPT outperforms general-purpose models including GPT-4, Llama-3, and Mistral in several telecom-domain tasks.
@@ 5.2 The RAG Results
Our experiments presented in Fig. 4 demonstrate that integrating RAG significantly improves model accuracy across all tested LLMs. Notably, GPT-4o hurst2024gpt exhibited a +11.9% improvement, suggesting that even sophisticated commercial models significantly benefit from structured retrieval of regulatory contexts. DeepSeek-R1-Distill-Qwen-14B guo2025deepseek exhibited the largest absolute improvement, +23%. Smaller models like DeepSeek-R1-Distill-Qwen-1.5B guo2025deepseek showed modest but notable gains of +3%, indicating that retrieval contexts help models of all scales, albeit differently. The figure 5 shows an example of how RAG efficiently helped GPT-4o to get the correct answer. More examples are provided in the Appendix, Section B.
GPT-4o with RAG context Answer: 174-174 dB(W/m2) in any 4kHz band Explanation: From the provided context: ”…the maximum power flux-density produced at the geostationary-satellite orbit by any Earth exploration-satellite service space station shall not exceed 174-174 dB(W/m2) in any 4 kHz band.”
This directly matches the value. The negative sign is correct—PFD values are often expressed in negative dB(W/m2) because they are very small.
@@ Appendix A Additional Results
Appendix B Question-Answer Examples

### v2

@@ Abstract
We study question answering in the domain of radio regulations, a legally sensitive and high-stakes area. We propose a telecom-specific Retrieval-Augmented Generation (RAG) pipeline and introduce, to our knowledge, the first multiple-choice evaluation set for this domain, constructed from authoritative sources using automated filtering and human validation. To assess retrieval quality, we define a domain-specific retrieval metric, under which our retriever achieves approximately 97% accuracy. Beyond retrieval, our approach consistently improves generation accuracy across all tested models. In particular, while naively inserting documents without structured retrieval yields only marginal gains for GPT-4o (less than 1%), applying our pipeline results in nearly a 12% relative improvement. These findings demonstrate that carefully targeted grounding provides a simple yet strong baseline and an effective domain-specific solution for regulatory question answering. All code and evaluation scripts, along with our derived question–answer dataset, are available at https://github.com/Zakaria010/Radio-RAG.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language processing, achieving state-of-the-art performance in summarization, translation, and question answering. However, despite their versatility, LLMs are prone to generating false or misleading content, a phenomenon commonly referred to as hallucination fourati2025coherence; Huang_2025; sahoo2024comprehensivesurveyhallucinationlarge. While sometimes harmless in casual applications, such inaccuracies pose significant risks in domains that demand strict factual correctness, including medicine, law, and telecommunications. In these settings, misinformation can have severe consequences, ranging from financial losses to safety hazards and legal disputes.
@@
To address these challenges, we introduce a domain-specialized Retrieval-Augmented Generation (RAG) pipeline for regulatory question answering. Our system leverages authoritative external resources to ground LLM outputs, thereby reducing hallucinations and improving reliability lewis2021retrieval; gupta2024comprehensive. To evaluate this approach, we construct a dedicated dataset of regulatory questions and answers derived from the ITU Radio Regulations ITURadioRegulations2024online, validated through both automated checks and human review. To make this system directly accessible to practitioners, we deploy it as a custom Generative Pre-trained Transformer (GPT) assistant, which integrates our RAG backend within a conversational interface.
Our contributions are fourfold:
@@
    We present an extensive empirical evaluation across multiple LLMs, showing that our pipeline consistently improves accuracy (e.g., +11.9% absolute for GPT-4o hurst2024gpt) and achieves up to 97% retrieval accuracy.
  4. 4.
    We deploy Radio Regulations GPT, a custom GPT front-end powered by our RAG backend, providing an interactive, reference-backed assistant for querying the ITU Radio Regulations. [^1] Link to Radio Regulations GPT:
    https://chatgpt.com/g/g-687bac267f508191a97daf466eccfa50-radio-regulations-gpt
@@ 2.1 LLMs
LLMs are transformer-based neural networks vaswani2023attentionneed; xiao2025foundationslargelanguagemodels trained on massive text corpora to acquire broad language understanding and reasoning capabilities. Recent models such as GPT-4 hurst2024gpt and GPT-5 demonstrate remarkable progress, yet still exhibit hallucinations and limited long-term memory fourati2025coherence.
LLMs performance can be further enhanced through several complementary strategies. Supervised fine-tuning (SFT) adapts pretrained models to specific domains or tasks by updating their parameters on curated labeled data raffel2023exploringlimitstransferlearning; hu2021loralowrankadaptationlarge; xu2023parameterefficientfinetuningmethodspretrained, while reinforcement learning from human feedback (RLHF) aligns their behavior with human preferences through reward-based optimization ouyang2022traininglanguagemodelsfollow. Both methods directly modify the model’s parametric memory, embedding new knowledge into its weights, yet this knowledge remains static and can quickly become obsolete, especially in rapidly evolving or highly regulated domains ovadia2024finetuningretrievalcomparingknowledge.
In contrast, prompt learning zhou2022large; kharrat2024acing and prompt engineering sahoo2024systematic guide model behavior without altering its parameters, using carefully crafted instructions or few-shot exemplars to induce desired task behavior at inference time. Relatedly, chain-of-thought (CoT) wei2022chain prompting encourages models to generate intermediate reasoning steps before producing an answer, which improves multi-step reasoning but does not expand the model’s knowledge base. While these approaches enhance task adaptability and reasoning ability, they cannot update or extend the underlying knowledge stored in the model’s parameters, leaving LLMs vulnerable to hallucinations fourati2025coherence, especially in specialized or fast-changing domains.
@@ 3 Related Works
The telecommunications domain poses distinctive challenges for language models due to its dense technical standards, highly structured documents, and precise terminology requirements. Recent efforts have begun adapting Retrieval-Augmented Generation (RAG) frameworks to this space. Bornea et al. bornea2024telcorag introduce Telco-RAG, a system designed to process 3GPP specifications, while Saraiva et al. saraiva2024telcodpr propose Telco-DPR, which evaluates retrieval models on a hybrid dataset combining textual and tabular inputs from telecom standards. Maatouk et al. maatouk2023teleqna contribute TeleQnA, a benchmark for assessing LLM knowledge of telecommunications. Subsequently Tele-LLMs maatouk2025telellmsseriesspecializedlarge were presented as a family of domain-specialized models trained on curated datasets (Tele-Data) and evaluated on a large-scale benchmark (Tele-Eval), demonstrating that targeted pretraining and parameter-efficient adaptation yield substantial improvements over general-purpose LLMs. More recently, Zou et al. zou2024telecomgpt developed TelecomGPT, a telecom‐specific LLM built via continual pretraining, instruction tuning, and alignment tuning, and evaluated on new benchmarks such as Telecom Math Modeling, Telecom Open QnA, and Telecom Code Tasks; TelecomGPT outperforms general-purpose models including GPT-4, Llama-3, and Mistral in several telecom‐domain tasks.
@@ 5.2 The RAG Results
Our experiments presented in Fig. 4 demonstrate that integrating RAG significantly improves model accuracy across all tested LLMs. Notably, GPT-4o hurst2024gpt exhibited a +11.9% improvement, suggesting that even sophisticated commercial models significantly benefit from structured retrieval of regulatory contexts. DeepSeek-R1-Distill-Qwen-14B guo2025deepseek exhibited the largest absolute improvement, +23%. Smaller models like DeepSeek-R1-Distill-Qwen-1.5B guo2025deepseek showed modest but notable gains of +3%, indicating that retrieval contexts help models of all scales, albeit differently. Examples are provided in the Appendix, Section C.
@@ 6 Discussion
Finally, to illustrate the practicality of our approach, we deployed the pipeline as a custom GPT, named Radio Regulations GPT, integrated with ChatGPT via an Actions endpoint. This deployment demonstrates that our RAG architecture can serve as a reliable conversational assistant for practitioners, maintaining grounding in the official ITU corpus while remaining easily updatable as regulations evolve. For more details refer to Appendix B.
@@ Appendix A Additional Results
Appendix B Deployment as Radio Regulations GPT
To make our system easily accessible to practitioners and regulators, we deploy the proposed RAG pipeline as a custom GPT, denoted Radio Regulations GPT. The RAG backend is exposed as a FastAPI service with a simple retrieval endpoint described via an OpenAPI schema:
POST /query: takes a JSON payload with the user question (query) and a fixed number of retrieved chunks (top_k), and returns the most relevant snippets from the indexed ITU Radio Regulations.
This API is integrated into ChatGPT through the Actions interface, so that the Custom GPT can call our backend directly. We configure Radio Regulations GPT with explicit system instructions that: (i) detect when a user request concerns the ITU Radio Regulations or related spectrum-policy questions; (ii) in such cases, call the /query endpoint with the full user message as input; (iii) treat the returned snippets as the primary evidence for the answer and synthesize a concise response grounded in these passages; (iv) include precise references to the relevant RR articles, numbers, and tables; and (v) fall back to standard LLM behavior without calling the API for clearly out-of-scope queries. This deployment ensures that answers remain faithful to the official ITU text while providing a convenient conversational interface.
Appendix C Question-Answer Examples
The figure 5 shows an example of how RAG efficiently helped GPT-4o [hurst2024gpt] to get the correct answer.
GPT-4o with RAG context Answer: 174-174 dB(W/m2) in any 4kHz band Explanation: From the provided context: ”…the maximum power flux-density produced at the geostationary-satellite orbit by any Earth exploration-satellite service space station shall not exceed 174-174 dB(W/m2) in any 4 kHz band.” This directly matches the value. The negative sign is correct—PFD values are often expressed in negative dB(W/m2) because they are very small.